A crate of oranges was shared equally among 15 people. 5 people gave up 3/5 of their share. As a result, the remaning people receivedd 12 more oranges each. How many oranges were there in the crate at first?
1
Expert's answer
2012-11-21T08:17:49-0500

2/5*5x = 120
2x = 120
x =60
Crate of oranges include 15x, so the total quantity of oranges = 60*15 = 900
Answer : 900